# Affine Symmetry and the Group-Theoretic Basis of the Unruh Effect

### v1

@@ I Introduction
The Unruh effect stands as one of the most striking manifestations of the interplay between quantum field theory and the geometry of spacetime. It reveals that the notion of a particle is inherently observer-dependent: a uniformly accelerated observer in Minkowski spacetime perceives the inertial vacuum as a thermal state with temperature proportional to the norm of its acceleration Unruh. The existence of vacuum states that appear “empty” only to specific classes of observers is a characteristic feature of quantum fields in non-inertial frames and in general curved spacetimes. This stems from the fact that different observers can adopt different definitions of positive frequency for the field modes, associated to different choices of time-like Killing vectors to generate translations in time, leading to distinct constructions of the one-particle Hilbert space and the associated Fock space Fulling:1972md; Ashtekar:1975zn. The celebrated result by Hawking, that black holes radiate thermally with a temperature inversely proportional to their mass Hawking_1, and the fact that cosmological horizons also carry an associated temperature Gibbons:1977mu are both manifestations of the same underlying phenomenon. Beyond its significance in horizon thermodynamics, the Unruh temperature of local Rindler horizons plays a central role in Jacobson’s suggestive derivation of Einstein equations from a Clausius relation linking horizon entropy to the energy flux through the horizon Jacobson:1995ab. This perspective remains highly relevant in current debates exploring the connection between horizon entropy and space-time dynamics which rely on the same conceptual ingredient: the thermal behavior associated with causal horizons Jacobson:2015hqa; Svesko:2018qim; Alonso-Serrano:2020pcz; Dorau:2025hmq.
@@
A wide range of derivations of the Unruh effect exist, each offering a different balance of rigor, simplicity, and physical intuition. Canonical, textbook treatments based on Bogoliubov transformations between Minkowski and Rindler modes clearly expose the origin of the effect due to the exisentce of inequivalent vacuum states but are technically demanding and and offer limited immediate physical intuition.Unruh; Crisp_Higu. Detector-based derivations provide an operational interpretation through the excitation spectrum of an Unruh–DeWitt detector, yet depend on idealized pointlike couplings, switching protocols, and subtle regularization schemes Donoghue:1983; Takagi:1986kn; Louko:2006zv. Euclidean approaches link thermal behavior to periodic imaginary time or loop contributions, but rest on analytic continuation Christensen:1978tw; Rajeev:2019bzv while more algebraic formulations reveal the general KMS structure associated with bifurcate Killing horizons, though at the cost of mathematical abstraction Sewell:1982zz; Kay:1988mu; Sorce:2024zme. Other semiclassical, tunneling arguments reproduce the Unruh temperature but typically rely on near-horizon WKB approximations and heuristic identifications of “inside” and “outside” degrees of freedom Kerner:2006vu.
@@ II Light-cone fields and the Unruh effect in two spacetime dimensions
The components of the decomposition (3) are independent since eq. (1) in light-cone coordinates reads $\partial_{u}\partial_{v}\phi(u,v)=0$ libro_wald whose general solution can be written as the sum of an arbitrary function of $u$ and an arbitrary function of $v$. Because these two sectors do not couple in the equation of motion, the theory exhibits a chiral structure with $\chi(u)$ and $\psi(v)$ evolving independently. Using the Heaviside step-function, the solutions $\psi(v)$ and $\chi(u)$ can be written as
@@
Uniformly accelerated observers follow hyperbolic trajectories in Minkowski spacetime (see Arzano:2024ogp for a pedagogic discussion), which can be described by Rindler coordinates $(\tau,\rho)$, defined in the right (R-) Rindler wedge $|t|<x$ by
@@
where $B(\omega)$ and $B^{\dagger}(\omega)$ are, again, expansion coefficients which we write as annihilation and creation operators for an uniformly accelerated observer, respectively. In order to obtain an expression of $B(\omega)$ and $B^{\dagger}(\omega)$ in terms of the annihilation and creation operators of an inertial observer $a(k)$ and $a^{\dagger}(k)$ we make use of the Mellin transform of $\psi_{+}(v)$ (see eg. Volov), defined by
@@
The derivation of the Unruh spectrum above may at first appear to be based on a technical shortcut obtained by applying the Mellin transform to quantum fields on the light cone. However, as we show in the following sections, the emergence of the Mellin transform reflects a deeper group-theoretical structure of the Unruh effect, rooted in the role of affine transformations acting on the light cones.
@@ III Light-cone evolution and the $ax+b$ group
showing a decomposition of the two-dimensional Weyl-Poincaré algebra in two identical sub-algebras Arzano:2018oby of the form
@@
There exist two nontrivial unitary irreducible representations of the affine group paper_rappr; libr_rapp that can be realized on the Hilbert spaces of functions $\mathcal{H}^{\pm}=\mathcal{L}^{2}\left(\mathbb{R}^{\pm},\frac{dk}{k}\right)$, whose elements can be denoted by kets $|k\rangle_{\pm}$ with invariant inner products
@@
showing that dilations move the states along the momentum axis, but they do not mix positive and negative sectors. Thus, since translations form a normal abelian subgroup, each ket $|k\rangle_{\pm}$ defines a character that is carried into another character under dilations, so one obtains full irreducible unitary representations without ever mixing the positive and negative momentum sectors. This construction is a special case of Mackey’s theory of induced representations raczka1986theory. We can write the actions (26) in terms of generators, which we denote with $P$ and $R$, eqs. (26) are given by
@@ IV Inequivalent Minkowski and Rindler modes in one-dimensional affine group representations
We will now consider the representations of the affine group above as the Hilbert spaces of a quantum system. Such system can be identified with a 0+1-dimensional quantum field theory or as the states of a one-dimensional quantum system with affine symmetry proc_2; proc_1. Since the unitary transformations (26) can be seen as the light-cone counterpart of Poincaré transformations, the states $|k\rangle_{+}$, eigenstates of the generator of translations, $P$ can be seen as one-particle states which we dub “P-particle states”. In terms of such P-eigenstates the $0+1$-dimensional affine field $\psi(t)=\langle t|\psi\rangle$ can be expanded as MosesQuesada1972ScaleEuclidean
@@
defining a reducible representation of the affine group on the real line. Indeed, irreducible representations on the whole line $t\in\mathbb{R}$ are defined considering only positive or negative frequency $k$. Moreover, we observe that the bra-ket
@@
From (37) and (38) it is clear that time evolution, generated by $P$, covers the whole real line while the action of the generator $R$ covers the half-line.
@@
To see that this is indeed the case, let us introduce a new coordinate $\xi$ “adapted” to dilations i.e. such that $D(\lambda)$ acts on it as a ordinary translation. Considering (38), it is clear that the change of variable is
@@
$|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int^{+\infty}_{0}\frac{dk}{k}k^{i\omega}|k\rangle_{+}\,$, (46)
@@
$\langle\omega|k\rangle_{+}=\frac{1}{\sqrt{2\pi}}k^{-i\omega}.$ (47)
@@
$R\,|\omega\rangle=\omega|\omega\rangle,\qquad P\,|\omega\rangle=|\omega-i\rangle\,$, (48)
@@
for translations we have libr_rapp
$T(\beta)\,|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int^{+\infty}_{0}\frac{dk}{k}\ k^{i\omega}e^{-i\beta k}|k\rangle_{+}$ (50)
@@
$|k\rangle_{+}=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}d\omega^{\prime}\;k^{-i\omega^{\prime}}\,|\omega^{\prime}\rangle,$ (51)
@@
$T(\beta)|\omega\rangle=\int_{-\infty}^{\infty}d\omega^{\prime}\;\left[\frac{1}{2\pi}\int_{0}^{\infty}\frac{dk}{k}\;k^{\,i(\omega-\omega^{\prime})}\,e^{-i\beta k}\right]\,|\omega^{\prime}\rangle.$ (52)
@@
$K(\omega,\omega^{\prime};\beta)=\frac{1}{2\pi}\int_{0}^{\infty}\frac{dk}{k}\;k^{\,i(\omega-\omega^{\prime})}\,e^{-i\beta k}.$ (53)
@@
and analytically continuing to $a=i\beta$, $s=i(\omega-\omega^{\prime})$, we find
$K(\omega,\omega^{\prime};\beta)=\frac{1}{2\pi}\,(i\beta)^{-\,i(\omega-\omega^{\prime})}\,\Gamma\!\big(i(\omega-\omega^{\prime})\big).$ (55)
@@
$\psi(t)=\frac{1}{2\pi}\int^{+\infty}_{-\infty}\int^{+\infty}_{0}\frac{dk}{k}(e^{ikt}k^{i\omega}\langle\omega|\psi\rangle+e^{-ikt}k^{-i\omega}\langle\omega|\psi\rangle^{*}),$ (57)
@@
We notice that the integral in eq.(57) is not defined for $\omega\in\mathbb{R}$, hence we regularize it by promoting $\omega$ to a complex variable considering $\omega\rightarrow\omega-i\lambda$ with $\lambda\in(0,1)$.
$\psi(t)=\frac{1}{2\pi}\int^{+\infty}_{0}\frac{dk}{k}\int^{+\infty}_{-\infty}d\omega(e^{ikt}k^{\lambda+i\omega}\ \langle\omega|\psi\rangle+e^{-ikt}k^{\lambda-i\omega}\ \langle\omega|\psi\rangle^{*}).$ (59)
@@
$\psi(t)$ $=\Theta(t)\frac{1}{2\pi}\int^{+\infty}_{-\infty}d\omega\biggl[\Gamma(\lambda+i\omega)t^{-\lambda-i\omega}e^{i\frac{\pi}{2}(+\lambda+i\omega)}\ \langle\omega|\psi\rangle+\Gamma(\lambda-i\omega)t^{-\lambda+i\omega}e^{-i\frac{\pi}{2}(+\lambda-i\omega)}\ \langle\omega|\psi\rangle^{*}\biggr]$ (60)
$+\Theta(-t)\frac{1}{2\pi}\int^{+\infty}_{-\infty}d\omega\biggl[\Gamma(\lambda+i\omega)(-t)^{-\lambda-i\omega}e^{-i\frac{\pi}{2}(+\lambda+i\omega)}\ \langle\omega|\psi\rangle+\Gamma(\lambda-i\omega)(-t)^{-\lambda+i\omega}e^{+i\frac{\pi}{2}(+\lambda-i\omega)}\ \langle\omega|\psi\rangle^{*}\biggr].$
@@
$\psi(t)$ $=\Theta(t)\Biggl[\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ t^{-\lambda}\Gamma(\lambda+i\omega)t^{-i\omega}[e^{+i\frac{\pi}{2}\lambda-\frac{\pi}{2}\omega}\ b_{+}(\omega)+e^{-i\frac{\pi}{2}\lambda+\frac{\pi}{2}\omega}\ b_{-}^{*}(\omega)]$ (61)
$+\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ t^{-\lambda}\Gamma(\lambda-i\omega)t^{i\omega}[e^{-i\frac{\pi}{2}\lambda-\frac{\pi}{2}\omega}\ b_{+}^{*}(\omega)+e^{+i\frac{\pi}{2}\lambda+\frac{\pi}{2}\omega}\ b_{-}(\omega)]\Biggr]$
$+\Theta(-t)\Biggl[\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ (-t)^{-\lambda}\Gamma(\lambda+i\omega)(-t)^{-i\omega}[e^{-i\frac{\pi}{2}\lambda+\frac{\pi}{2}\omega}\ b_{+}(\omega)+e^{+i\frac{\pi}{2}\lambda-\frac{\pi}{2}\omega}\ b_{-}^{*}(\omega)]$
$+\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ (-t)^{-\lambda}\Gamma(\lambda-i\omega)(-t)^{i\omega}[e^{+i\frac{\pi}{2}\lambda+\frac{\pi}{2}\omega}\ b_{+}^{*}(\omega)+e^{-i\frac{\pi}{2}\lambda-\frac{\pi}{2}\omega}\ b_{-}(\omega)]\Biggr],$

### v2

@@ I Introduction
The Unruh effect stands as one of the most striking manifestations of the interplay between quantum field theory and the geometry of spacetime. It reveals that the notion of a particle is inherently observer-dependent: a uniformly accelerated observer in Minkowski spacetime perceives the inertial vacuum as a thermal state with temperature proportional to the norm of its acceleration Unruh (1976). The existence of vacuum states that appear “empty” only to specific classes of observers is a characteristic feature of quantum fields in non-inertial frames and in general curved spacetimes. This stems from the fact that different observers can adopt different definitions of positive frequency for the field modes, associated to different choices of time-like Killing vectors to generate translations in time, leading to distinct constructions of the one-particle Hilbert space and the associated Fock space Fulling (1973); Ashtekar and Magnon (1975). The celebrated result by Hawking, that black holes radiate thermally with a temperature inversely proportional to their mass Hawking (1975), and the fact that cosmological horizons also carry an associated temperature Gibbons and Hawking (1977) are both manifestations of the same underlying phenomenon. Beyond its significance in horizon thermodynamics, the Unruh temperature of local Rindler horizons plays a central role in Jacobson’s suggestive derivation of Einstein equations from a Clausius relation linking horizon entropy to the energy flux through the horizon Jacobson (1995). This perspective remains highly relevant in current debates exploring the connection between horizon entropy and space-time dynamics which rely on the same conceptual ingredient: the thermal behavior associated with causal horizons Jacobson (2016); Svesko (2019); Alonso-Serrano and Liška (2020); Dorau and Much (2025).
@@
A wide range of derivations of the Unruh effect exist, each offering a different balance of rigor, simplicity, and physical intuition. Canonical, textbook treatments based on Bogoliubov transformations between Minkowski and Rindler modes clearly expose the origin of the effect due to the exisentce of inequivalent vacuum states but are technically rather involved and offer limited immediate physical intuition.Unruh (1976); Crispino et al. (2008). Detector-based derivations provide an operational interpretation through the excitation spectrum of an Unruh–DeWitt detector, yet depend on idealized pointlike couplings, switching protocols, and subtle regularization schemes Donoghue and Holstein (1984); Takagi (1986); Louko and Satz (2006). Euclidean approaches link thermal behavior to periodic imaginary time or loop contributions, but rest on analytic continuation Christensen and Duff (1978); Rajeev and Padmanabhan (2020) while more algebraic formulations reveal the general KMS structure associated with bifurcate Killing horizons, though at the cost of mathematical abstraction Sewell (1982); Kay and Wald (1991); Sorce (2024). Other semiclassical, tunneling arguments reproduce the Unruh temperature but typically rely on near-horizon WKB approximations and heuristic identifications of “inside” and “outside” degrees of freedom Kerner and Mann (2006).
@@ II Light-cone fields and the Unruh effect in two spacetime dimensions
The components of the decomposition (3) are independent since eq. (1) in light-cone coordinates reads $\partial_{u}\partial_{v}\phi(u,v)=0$ Wald (1995) whose general solution can be written as the sum of an arbitrary function of $u$ and an arbitrary function of $v$. Because these two sectors do not couple in the equation of motion, the theory exhibits a chiral structure with $\chi(u)$ and $\psi(v)$ evolving independently. Using the Heaviside step-function, the solutions $\psi(v)$ and $\chi(u)$ can be written as
@@
Uniformly accelerated observers follow hyperbolic trajectories in Minkowski spacetime (see Arzano et al. (2024) for a pedagogic discussion), which can be described by Rindler coordinates $(\eta,\chi)$, defined in the right (R-) Rindler wedge $|t|<x$ by
@@
where $B(\omega)$ and $B^{\dagger}(\omega)$ are, again, expansion coefficients which we write as annihilation and creation operators for an uniformly accelerated observer, respectively. In order to obtain an expression of $B(\omega)$ and $B^{\dagger}(\omega)$ in terms of the annihilation and creation operators of an inertial observer $a(k)$ and $a^{\dagger}(k)$ we make use of the Mellin transform of $\psi_{+}(v)$ (see eg. Aref’eva and Volovich (2013)), defined by
@@
The derivation of the Unruh spectrum above may at first appear to be based on a technical shortcut consisting in applying the Mellin transform to quantum fields on the light cone. However, as we show in the following sections, the emergence of the Mellin transform reflects a deeper group-theoretical structure of the Unruh effect, rooted in the role of affine transformations acting on the light cones.
@@ III Light-cone evolution and the $ax+b$ group
showing a decomposition of the two-dimensional Weyl-Poincaré algebra in two identical sub-algebras Arzano and Kowalski-Glikman (2019b) of the form
@@
There exist two nontrivial unitary irreducible representations of the affine group Zeitlin (2016); Vilenkin et al. (1968) that can be realized on the Hilbert spaces of functions $\mathcal{H}^{\pm}=\mathcal{L}^{2}\left(\mathbb{R}^{\pm},\frac{dk}{k}\right)$, whose elements can be denoted by kets $|k\rangle_{\pm}$ with invariant inner products
@@
showing that dilations move the states along the momentum axis, but they do not mix positive and negative sectors. Thus, since translations form a normal abelian subgroup, each ket $|k\rangle_{\pm}$ defines a character that is carried into another character under dilations, so one obtains full irreducible unitary representations without ever mixing the positive and negative momentum sectors. This construction is a special case of Mackey’s theory of induced representations Raczka and Barut (1986). We can write the actions (26) in terms of generators, which we denote with $P$ and $R$, eqs. (26) are given by
@@ IV Inequivalent Minkowski and Rindler modes in one-dimensional affine group representations
We will now consider the representations of the affine group above as the Hilbert spaces of a quantum system. Such system can be identified with a 0+1-dimensional quantum field theory or as the states of a one-dimensional quantum system with affine symmetry Arzano and Kowalski-Glikman (2019a); Arzano (2020). Since the unitary transformations (26) can be seen as the light-cone counterpart of Poincaré transformations, the states $|k\rangle_{+}$, eigenstates of the generator of translations, $P$ can be seen as one-particle states which we dub “P-particle states”. In terms of such P-eigenstates the $0+1$-dimensional affine field $\psi(t)=\langle t|\psi\rangle$ can be expanded as Moses and Quesada (1972)
@@
defining a reducible representation of the affine group on the real line. Indeed, irreducible representations on the whole line $t\in\mathbb{R}$ are defined considering only positive or negative frequency $k$. Moreover, we observe that the bra-kets
@@
From (37) and (38) it is clear that time evolution, generated by $P$, covers the whole real line while transformations generated by $R$ cover only the positive or the negative half-lines.
@@
To see that this is indeed the case, let us introduce a new coordinate $\xi$ “adapted” to dilations i.e. such that $D(\lambda)$ acts on it as a ordinary translation. Considering (38), it is clear that such change of variable is given by
@@
$|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int^{+\infty}_{0}\frac{dk}{k}k^{-i\omega}|k\rangle_{+}\,$, (46)
@@
$\langle\omega|k\rangle_{+}=\frac{1}{\sqrt{2\pi}}\,k^{i\omega}.$ (47)
@@
$R\,|\omega\rangle=\omega|\omega\rangle,\qquad P\,|\omega\rangle=|\omega{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\rangle\,$, (48)
@@
for translations we have Vilenkin et al. (1968)
$T(\beta)\,|\omega\rangle=\frac{1}{\sqrt{2\pi}}\int^{+\infty}_{0}\frac{dk}{k}\ k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}e^{-i\beta k}|k\rangle_{+}$ (50)
@@
$|k\rangle_{+}=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}d\omega^{\prime}\;k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega^{\prime}}\,|\omega^{\prime}\rangle,$ (51)
@@
$T(\beta)|\omega\rangle=\int_{-\infty}^{\infty}d\omega^{\prime}\;\left[\frac{1}{2\pi}\int_{0}^{\infty}\frac{dk}{k}\;k^{\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i(\omega-\omega^{\prime})}\,e^{-i\beta k}\right]\,|\omega^{\prime}\rangle.$ (52)
@@
$K(\omega,\omega^{\prime};\beta)=\frac{1}{2\pi}\int_{0}^{\infty}\frac{dk}{k}\;k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\,i(\omega-\omega^{\prime})}\,e^{-i\beta k}.$ (53)
@@
and analytically continuing to $a=i\beta$, $s={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i(\omega-\omega^{\prime})$, we find
$K(\omega,\omega^{\prime};\beta)=\frac{1}{2\pi}\,(i\beta)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\,i(\omega-\omega^{\prime})}\,\Gamma\!\big({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i(\omega-\omega^{\prime})\big).$ (55)
@@
$\psi(t)=\frac{1}{2\pi}\int^{+\infty}_{-\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d\omega}\int^{+\infty}_{0}\frac{dk}{k}(e^{ikt}k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}\langle\omega|\psi\rangle+e^{-ikt}k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega}\langle\omega|\psi\rangle^{*}),$ (57)
@@
We notice that the integral in eq.(57) is not defined for $\omega\in\mathbb{R}$, hence we regularize it by promoting $\omega$ to a complex variable considering $\omega\rightarrow\omega{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\lambda$ with $\lambda\in(0,1)$.
$\psi(t)=\frac{1}{2\pi}\int^{+\infty}_{0}\frac{dk}{k}\int^{+\infty}_{-\infty}d\omega(e^{ikt}k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}\ \langle\omega|\psi\rangle+e^{-ikt}k^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega}\ \langle\omega|\psi\rangle^{*}).$ (59)
@@
[…]
@@
$\psi(t)$ $=\Theta(t)\Biggl[\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\lambda}\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega)t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega}[e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{\pi}{2}\omega}\ b_{+}(\omega)+e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{\pi}{2}\omega}\ b_{-}^{*}(\omega)]$ (61)
$+\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\lambda}\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega)t^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}[e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{\pi}{2}\omega}\ b_{+}^{*}(\omega)+e^{+i\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{\pi}{2}\omega}\ b_{-}(\omega)]\Biggr]$
$+\Theta(-t)\Biggl[\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ (-t)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\lambda}\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega)(-t)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega}[e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{\pi}{2}\omega}\ b_{+}(\omega)+e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{\pi}{2}\omega}\ b_{-}^{*}(\omega)]$
$+\int^{+\infty}_{0}\frac{d\omega}{2\pi}\ (-t)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\lambda}\Gamma({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\omega)(-t)^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\omega}[e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\frac{\pi}{2}\omega}\ b_{+}^{*}(\omega)+e^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}i\frac{\pi}{2}\lambda{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}\frac{\pi}{2}\omega}\ b_{-}(\omega)]\Biggr],$